solve by the subst
Solve by the substitution method.
9x+3y=-30
-3x+y=20
What is the solution of the system?
The solution is____?
(Type an ordered pair)

Answer by mananth(16946)

9x+3y=-30 .............1
y=20+3x ---------2
/1=
y=20+3x
Plug the value of y in (1)
9x+3(20+3x)=-30
9x+60+9x=-30
9x+9x= -30 -60
18x=-90
/18
x=-5
Plug the value of x in (1)
9x+3y=-30
9*-5+3y=-30
-45+3y=-30
3y=15
/3
y=5
(-5,5)